A horizontal rectangular surface has dimensions 2.55 cm by 3.15 cm and is in a uniform magnetic field that is directed at an angle of 32.5 above the horizontal. 


Part A 

What must the magnitude of the magnetic field be in order to produce a flux of 4.10x10-4 Wb through the surface? 
Answer #1

given that

area of rectangular surface A = 2.55*3.15*10^-4 = 8.0325*10^-4 m^2

angle = theta = 32.5 degrees

magnetic flux = 4.1*10^-4 Wb

from the relation

magnetic flux = BAcos(90-theta)

B = 4.1*10^-4/8.0325*10^-4*cos(90-32.5) = 0.951 T